# Position: On-Premises LLM Deployment Demands a Middle Path: Preserving Privacy Without Sacrificing Model Confidentiality

### v2

@@ Abstract
Current LLM customization typically relies on two deployment strategies: closed-source APIs, which require users to upload private data to external servers, and open-weight models, which allow local fine-tuning but pose misuse risks. In this position paper, we argue that (1) deploying closed-source LLMs within user-controlled infrastructure (on-premises deployment) enhances data privacy and mitigates misuse risks, and (2) a well-designed on-premises deployment must ensure model confidentiality—by preventing model theft—and offer privacy-preserving customization. Prior research on small models has explored securing only the output layer within hardware-secured devices to balance confidentiality and customization efficiency. However, we show that this approach is insufficient for defending large-scale LLMs against distillation attacks. We therefore introduce a semi-open deployment framework that secures only a few, carefully chosen layers, achieving distillation resistance comparable to fully secured models while preserving fine-tuning flexibility. Through extensive experiments, we show that securing bottom layers significantly reduces functional extraction risks. Our findings demonstrate that privacy and confidentiality can coexist, paving the way for secure on-premises AI deployment that balances usability and protection.
Machine Learning, Secure on-premises LLM Deployment
@@ 1 Introduction
Vendors of Large Language Models (LLMs) have introduced advanced models with remarkable capabilities to address diverse user needs (Minaee et al., 2024; Zhao et al., 2023). To enable customization and drive industry progress, vendors typically adopt two approaches. Closed-source vendors, like OpenAI, provide fine-tuning APIs that allow users to upload data to customize proprietary models such as GPT-4. In contrast, companies like Meta offer open-weight models, such as Llama3 (Dubey et al., 2024), which users can adapt within their own infrastructure, ensuring greater flexibility and control.
However, both approaches present significant limitations for privacy-sensitive users, such as financial institutions, healthcare organizations, and government agencies, which prioritize data security. Fine-tuning APIs from closed-source vendors offer encryption for sensitive data and comply with strict privacy regulations (Dong et al., 2023; Pang et al., 2024). However, their security fundamentally depends on vendor trust, raising ethical concerns and exposing users to potential data breaches (Sun et al., 2023).
Alternatively, fine-tuning open-weight models within private infrastructure—commonly referred to as on-premises deployment, where all data and model adaptation occur locally—ensures data protection and customization (Nevo et al., 2024). However, full disclosure of model architectures and weights heightens the risk of misuse by malicious actors, enabling misinformation generation, automated cyberattacks, and security bypasses (Hendrycks et al., 2023). These risks discourage vendors from releasing state-of-the-art (SOTA) models as open-weight, as uncontrolled distribution could lead to widespread harm. Beyond security concerns, maintaining high-quality open-weight models imposes substantial computational and financial costs (Wolfe et al., 2024), further disincentivizing full disclosure.
In this position paper, we argue that deploying closed-source SOTA models on-premises with restricted access to authorized users offers a middle path to mitigate privacy leakage and model misuse risks. However, prior research has shown that hardware-based extraction attacks can recover model parameters and architectures directly from CPUs and memory (Hu et al., 2020) within local infrastructures, enabling unauthorized users to steal proprietary model information from the deployment server. This underscores the need for stronger security measures. Therefore, we further argue that a well-designed on-premises deployment strategy must fully prevent model theft while preserving data privacy, ensuring no reliance on untrusted servers.
Our work provides an affirmative step toward addressing the fundamental challenge of balancing security and customization in on-premises LLM deployment. Existing approaches have sought to protect proprietary models using Trusted Execution Environments (TEEs) (Narra et al., 2019), but fully enclosing models within TEEs incurs prohibitive computational costs, limiting their practicality (Nayan et al., 2024). Prior research has attempted to mitigate this trade-off by securing only critical layers, such as the output layer, while leaving the remaining layers exposed for fine-tuning (Zhang et al., 2024a; Mo et al., 2020). However, our findings align with prior work showing that securing only the output layer is insufficient against distillation attacks (Zanella-Beguelin et al., 2021). Extending these attacks to Llama2-70B, we confirm that this vulnerability persists at scale, enabling near-complete functionality distillation in six domains, as shown in Figure 1. These vulnerabilities raise skepticism about whether model confidentiality and customization can truly coexist in on-premises deployment, highlighting the need for a security paradigm beyond final-layer protection. Without a viable solution, vendors risk intellectual property theft, while users must compromise data privacy when relying on external servers.
@@
Our work suggests that this dilemma can be resolved. We propose a theoretically inspired semi-open deployment strategy that secures only a few, carefully chosen layers, balancing security and customization while mitigating distillation risks. This enables secure on-premises AI adoption in privacy-sensitive industries such as healthcare, finance, and government. More broadly, it highlights the need for hybrid security solutions that balance usability and protection, lowering barriers to responsible AI deployment.
By demonstrating that confidentiality and customization can coexist, our work shifts the narrative from skepticism to opportunity, paving the way for AI-driven innovation in regulated sectors. Future research should explore optimal layer selection, adaptive security mechanisms, and distillation-resistant architectures to further enhance model confidentiality. We hope this work inspires the AI community to develop scalable solutions that ensure AI remains both secure and widely accessible. We summarize our contribution as follows.
  • Semi-open on-premises LLM deployment. We propose a semi-open deployment framework that secures a single, carefully chosen layer, achieving distillation resistance comparable to full encryption while preserving customization flexibility comparable to full parameter fine-tuning.
  • Layer-wise security transitions. We provide the first theoretical result identifying a transition layer that resists distillation attacks, revealing fundamental differences in security across model layers.
  • Comprehensive empirical evaluations. We have conducted extensive empirical evaluations across five models (Llama2 (70B,7B), Mistral-B, Phi-2 (2.7B), and Phi-1.5 (1.3B) ) under three distillation strategies with a 51k attack set, comparing its performance with three baselines (DarkneTZ, SAP-DP, Fully-secured).
  • Discussion of alternative positions. We objectively examine alternative positions, and highlight their benefits and limitations, reinforcing the need for a new framework that balances security, privacy, and customization. Our discussion aims to stimulate further advancements and foster new solutions in this evolving field.
2 Related Works
2.1 Data Privacy Risks
Using LLM services for customization introduces significant data privacy risks, as user data is exposed during transmission, storage, and processing (Li et al., 2024b). To address this, prior research has explored Homomorphic Encryption (Lee et al., 2022), which allows computations on encrypted data, and Differential Privacy (Wei et al., 2021), which injects noise to prevent sensitive data memorization. However, these methods suffer from high computational overhead (Zhou et al., 2024) and limited protection coverage (Wei et al., 2020). A more effective approach is on-premises deployment, which ensures data remains under user control (Nevo et al., 2024). While this mitigates privacy concerns, it shifts the risk to LLM vendors, who lose control over model usage and face heightened risks of theft and misuse (Hendrycks et al., 2023).
2.2 Model Theft Risks
Deploying LLMs on user-controlled infrastructure increases exposure to model extraction threats (Atli et al., 2020). Attackers can recover model parameters through hardware-based techniques, including side-channel attacks targeting GPUs (Nayan et al., 2024) and bus monitoring attacks (Hu et al., 2020). In addition to direct parameter extraction, distillation attacks allow adversaries to replicate model functionality using only black-box access, a process known as functional distillation (Nevo et al., 2024; Xu et al., 2024; Ezzeddine et al., 2024). While distillation attacks have been extensively studied in smaller models, such as CNNs (Orekondy et al., 2018), BERT (Sanh et al., 2020; Zanella-Beguelin et al., 2021), and ReLU-based models (Canales-Martínez et al., 2024; Jagielski et al., 2020), their effectiveness against large-scale LLMs remains an open question. Our work extends these attacks to Llama2-70B and demonstrates that securing only the output layer remains insufficient to prevent near-complete functionality replication.
2.3 Traditional Approaches
Balancing data privacy and model security in on-premises deployment requires novel solutions. One direction is privacy-preserving federated deployment, where a model is split between user-controlled infrastructure and vendor-managed servers (Shen et al., 2023; Huang et al., 2024). Users fine-tune local layers while the vendor controls deeper layers, with privacy-preserving techniques like differential privacy (Zhou et al., 2024) protecting sensitive data. However, industries with strict data regulations, such as banking, healthcare, and government, often require full on-premises deployment (Schillaci, 2024; Guerra-Manzanares et al., 2023), ensuring models remain within their infrastructure while allowing customization.
To secure models in on-premises settings, research has explored hardware-based protections such as Arm TrustZone (Pinto & Santos, 2019) and secure execution environments (Zhang et al., 2024b; Li et al., 2024a). These approaches isolate computations to prevent unauthorized extraction but impose high resource demands and restrict customization (Mo et al., 2020). A more flexible alternative is layer-wise security, where only critical layers are protected while others remain exposed (Lin et al., 2024; Chen et al., 2024; Zhang et al., 2024a). Studies suggest different strategies, including securing shallow layers (Elgamal & Nahrstedt, 2020), intermediate layers (Shen et al., 2022), or the output layer (Huang et al., 2024). However, most research has focused on smaller models, leaving the effectiveness of different security placements in LLMs unclear.
Our work investigates the layer-wise security in LLMs, showing that securing bottom layers better resists distillation. This suggests that protecting a single, well-chosen layer can defend against attacks while maximizing fine-tuning flexibility, enabling more secure on-premises deployment.
@@ 3.1 Security Threat: Model Distillation
Semi-open LLMs. Let $f:\mathbb{R}^{n\times d}\rightarrow\mathcal{Y}$ denote a victim model that processes the feature matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ to produce an output in set $\mathcal{Y}$. Modern LLMs typically adopt a multi-layer architecture to capture complex patterns in the input data. Specifically, $f(\mathbf{X};\bm{\theta})=\varphi_{L}\circ...\circ\varphi_{1}(\mathbf{X})$ is a composition of multiple decoder layers $\varphi_{1},...,\varphi_{L}$, each shares the same architecture but with distinct parameters. The parameters of all layers are denoted by the vector $\bm{\theta}$. We consider a semi-open setting, in line with (Zanella-Beguelin et al., 2021) and (Xu et al., 2021), where certain layers of the LLM are secured while others remain public.
@@
Adversary’s Objective. The adversary aims to replicate the functionality of a semi-open victim LLM, partially secured in a protected environment, by building a replacement model. The agreement between the victim and replicated models is measured through accuracy and fidelity on a specified testing dataset.
Adversary’s Knowledge. It is assumed that the adversary knows the architectures of both secured and unsecured modules, as prior work (Gou et al., 2021; Boix-Adsera, 2024) has shown that using the same architecture as the secured module significantly improves the effectiveness of distillation attacks. However, the adversary knows only the parameters of the unsecured module, while those of the secured module remain unknown.
@@ 3.2 Problem Formulation
What is the smallest secured set $\text{What is the smallest secured set }I\text{ such that }$
@@ 4.1 Security Transition in Deep Transformers
$\varphi_{i}(\mathbf{X};K_{i},Q_{i})=\mathbf{X}+\operatorname{softmax}\left(\frac{\mathbf{X}Q_{i}(\mathbf{X}K_{i})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)\mathbf{X}.$
@@
Assume that $\mathbb{P}_{\mathbf{X}\times Y}$ is defined on a countable domain $\mathcal{X}\times\mathcal{Y}$ with $\mathbf{0}_{n\times d}\notin\mathcal{X}$. Assume that parameter matrices $\{K_{i},Q_{i}\}_{i\geq 1}$ in the victim model $f$ have uniform bounded norms, i.e., $\|K_{i}\|\leq D$ and $\|Q_{i}\|\leq D$ for some $D>0$. There exists an $\alpha^{*}\in(0,1)$ depending on $D$ such that the following two statements are true.
(1) If $\alpha<\alpha^{*}$ and $\{K_{i},Q_{i}\}_{i\geq 1}$ are parameter matrices of the victim model, with $\hat{K}_{\alpha L}$ and $\hat{Q}_{\alpha L}$ as distilled parameters drawn from a continuous distribution on $\mathbb{R}^{n\times d}$, the matrix $\hat{f}_{\infty}(\mathbf{X})$ exists and almost surely has rank one for all $\mathbf{X}$.
(2) If $\alpha>\alpha^{*}$, there exists a victim model with parameter sequence $\{K_{i},Q_{i}\}_{i\geq 1}$ such that for any distilled parameters $\hat{K}_{\alpha L}$ and $\hat{Q}_{\alpha L}$, the matrix $\hat{f}_{\infty}(\mathbf{X})$ exists but has rank greater than one for some $\mathbf{X}$.
Remark 1: The proof is provided in Appendix A. This theorem demonstrates that if the distilled parameters of the bottom layers (i.e., $\alpha<\alpha^{*}$) are obtained through a randomized algorithm, such as stochastic gradient descent, with a continuous distribution supported on $\mathbb{R}^{n\times d}$, the distillation will certainly fail, as the feature matrix degenerate. In contrast, keeping the later layers secured (i.e., $\alpha>\alpha^{*}$) does not maintain this property, indicating that it is more effective to secure the bottom layers before the transition layer, rather than the later ones.
Theorem 1 shows that protecting bottom layers enhances resilience against model distillation attacks, indicating that securing layers starting from the first may be effective. In the next subsection, we propose SOLID, a preliminary LLM deployment framework solution that balance model protection and customization.
@@ 4.2 SOLID: Semi-Open Local Infrastructure Deployment Framework
We propose a method to approximately find the smallest bottom layer index set $I$ that satisfies $R(I)\leq(1+\varepsilon)R([L])$ for any small $\varepsilon>0$. A simple approach is to start with $I_{l}=\{1,\ldots,l\}$ for each $l$ beginning from 1, then evaluate the distillation ratio $R(I_{l})$ after the attack, and identify the smallest $l$ that meets the inequality. This extensive fine-tuning process is time-consuming, prompting the critical question: Can we create a fine-tuning-free metric that predicts LLM performance under model distillation attacks? Hence, our goal is to establish a metric directly correlated with the distillation ratio.
@@
Previous research (Choi et al., 2024; Bailly et al., 2022) indicates that the difference $\|\bm{\theta}_{\text{dist}}(I,\mathcal{D})-\bm{\theta}_{0}(I)\|_{2}$ is minimal in large networks compared to the dataset size $|\mathcal{D}|$. In models such as single-layer ReLU networks (Anthony et al., 1999; Zou et al., 2020), this difference scales as $\mathcal{O}\left(\frac{|\mathcal{D}|}{\sqrt{N}}\right)$ (Jacot et al., 2018; Wei et al., 2019), where $N$, the number of model parameters, far exceeds the dataset size in large language models (LLMs) (Dubey et al., 2024; Liu et al., 2024). The first term, independent of fine-tuning, dominates and effectively predicts the distillation ratio. We refer to this term as the Distillation Difficulty (DD$(I)$), defined as
@@
This score, which can be estimated using a sample average, represents the distilled model performance of the model when specific layers $I$ are secured. A higher DD$(I)$ suggests better security performance, indicating a lower distillation ratio $R(I)$. Therefore, our SOLID operates in the following way. SOLID begins by sampling evaluation data targeting general capabilities from the underlying distribution, and then computes DD$(I_{l})$ for each set of secured layers $I_{l}=\{1,...,l\}$ for $l=1,...,L$. SOLID stops at the smallest $l^{*}$ that satisfies DD$(I_{l^{*}})\geq(1-\varepsilon)\text{DD}([L])$.

### v3

# A Middle Path for On-Premises LLM Deployment: Preserving Privacy Without Sacrificing Model Confidentiality

@@ Abstract
Privacy-sensitive users require deploying large language models (LLMs) within their own infrastructure (on-premises) to safeguard private data and enable customization. However, vulnerabilities in local environments can lead to unauthorized access and potential model theft. To address this, prior research on small models has explored securing only the output layer within hardware-secured devices to balance model confidentiality and customization. Yet this approach fails to protect LLMs effectively. In this paper, we discover that (1) query-based distillation attacks targeting the secured top layer can produce a functionally equivalent replica of the victim model; (2) securing the same number of layers, bottom layers before a transition layer provide stronger protection against distillation attacks than top layers, with comparable effects on customization performance; and (3) the number of secured layers creates a trade-off between protection and customization flexibility. Based on these insights, we propose SOLID, a novel deployment framework that secures a few bottom layers in a secure environment and introduces an efficient metric to optimize the trade-off by determining the ideal number of hidden layers. Extensive experiments on five models (1.3B to 70B parameters) demonstrate that SOLID outperforms baselines, achieving a better balance between protection and downstream customization. Our code can be found at: https://github.com/OTTO-OTO/SOLID-OnPremiseDeployment.
A Middle Path for On-Premises LLM Deployment:
Preserving Privacy Without Sacrificing Model Confidentiality
Hanbo Huang1, Yihan Li2, Bowen Jiang1, Bo Jiang1, Lin Liu2, Ruoyu Sun3, Zhuotao Liu4, Shiyu Liang1 1Shanghai Jiao Tong University, 2National University of Defense Technology, 3Chinese University of Hong Kong (Shenzhen), 4Tsinghua University {hhuang417,lsy18602808513}@sjtu.edu.cn
@@ 1 Introduction
Vendors of Large Language Models (LLMs) have introduced advanced models with remarkable capabilities to address diverse user needs (Minaee et al., 2024; Zhao et al., 2023). To meet specific customization demands, vendors typically adopt two approaches. Closed-source vendors, such as OpenAI, provide fine-tuning APIs that allow users to upload data to customize proprietary models like GPT-4. In contrast, vendors like Meta offer open-weight models such as Llama3 (Dubey et al., 2024), which users can adapt within their own infrastructure, ensuring greater flexibility and control.
However, both approaches present notable limitations for privacy-sensitive users, such as healthcare organizations, who prioritize data security. Strict regulations prohibit these users from uploading sensitive data to third-party API services, necessitating on-premises deployment of LLMs, where data processing and model customization are confined to local infrastructure (Nevo et al., 2024). Although fine-tuning open-weight models in such environments offers a viable path for customization, full disclosure of model architectures and weights increases the risk of exploitation by malicious actors, who may circumvent safety mechanisms (Hendrycks et al., 2023). Consequently, vendors may be hesitant to release SOTA models as open-weight, since uncontrolled access could lead to significant harm. Moreover, maintaining high-quality open-weight models imposes considerable computational and financial burdens (Wolfe et al., 2024). These growing concerns highlight the importance of secure on-premises deployment of closed-source models, which preserves control and ensures regulatory compliance.
Despite the advantages, deploying closed-source LLMs locally introduces the risk of model theft. Unauthorized users can extract model parameters and architectures directly from CPUs and memory within local environments (Hu et al., 2020). To mitigate this, existing approaches use Trusted Execution Environments (TEEs) to protect proprietary models (Nayan et al., 2024; Narra et al., 2019). Yet, fully enclosing large models within TEEs results in prohibitive computational overhead, limiting their practicality (Li et al., 2024a).
Prior research has explored to mitigate this trade-off by securing only critical layers, such as the output layer, while leaving the remaining layers exposed for fine-tuning (Zhang et al., 2024b; Mo et al., 2020). However, studies have shown that even with only black-box access, adversaries may still be able to replicate the weights of DNNs (Tramèr et al., 2016; Truong et al., 2021). More recent works (Carlini et al., 2024; Finlayson et al., 2024) further suggests that the final-layer weights of large language models (LLMs) can be recovered from output logits alone, raising concerns about the robustness of such partial protection strategies. Consistent with previous findings, our results show that this partial protection remains vulnerable to distillation attacks (Zanella-Beguelin et al., 2021). When extended to Llama2-70B, we confirm that attackers can still extract nearly complete model functionality across six domains, as illustrated in Figure 1. These vulnerabilities raise skepticism about whether model confidentiality and customization can truly coexist in on-premises deployment, highlighting the need for security paradigms beyond output-layer protection.
In this paper, we show that this dilemma can be resolved. We begin by investigating the security-customization trade-off introduced by the placement of secured layers in LLMs. Specifically, we theoretically identify a transition layer in deep transformers, showing that securing bottom layers before this transition significantly reduces distillation success, while securing top layers has a more limited impact. Besides, we demonstrate that the number of secured layers creates a trade-off: securing more layers improves security but reduces customization flexibility. To optimize this trade-off, we introduce SOLID, a semi-open deployment framework that selectively secures a subset of bottom layers, using a distillation difficulty score to identify the optimal set for protection. Our experiments show that SOLID balances security and customization, achieving security comparable to fully secured models while maintaining strong customization flexibility, approaching full parameter fine-tuning. Our main contributions are as follows:
  • We extend query-based distillation attacks to LLMs, demonstrating that existing on-premises frameworks risk full functionality replication.
  • We identify the security-customization trade-off introduced by the placement of secured layers, and theoretically prove that securing bottom layers before the transition layer offers stronger protection with similar customization effects.
  • We discover that the number of secured layers affects both security and customization. We propose SOLID, which optimizes the security-customization trade-off by using a fine-tuning-free metric to secure minimal bottom decoder layers, protecting the model from distillation attacks while preserving customization flexibility.
  • We evaluate SOLID against three baselines across five models (1.3B to 70B parameters), assessing security across three distillation strategies on sixteen benchmarks and customization flexibility across six tasks. Extensive experiments show SOLID effectively balances security and customization, despite some limitations
@@ 2.1 Security Threat: Model Distillation
Adversary’s Objective. The adversary aims to replicate the functionality of a semi-open victim LLM, partially secured in a protected environment, by training a substitute model. This replica facilitates white-box analysis to identify vulnerabilities, enhancing black-box attacks on related model families (Sitawarin et al., 2024). The agreement between the victim and replica is assessed via accuracy and fidelity on a designated test set.
Adversary’s Knowledge. It is assumed that the adversary knows the architectures of both secured and unsecured modules, as prior work Gou et al. (2021); Boix-Adsera (2024) has shown that using the same architecture as the secured module significantly improves the effectiveness of distillation attacks. However, the adversary knows only the parameters of the unsecured module, while those of the secured module remain unknown.
@@ 2.2 Problem Formulation
[…]
@@ 3.1 Security Transition in Deep Transformers
$\varphi_{i}(\mathbf{X};K_{i},Q_{i})=\mathbf{X}+\operatorname{softmax}\!\left(\frac{\mathbf{X}Q_{i}(\mathbf{X}K_{i})^{\top}}{\sqrt{d_{Q}}\,\|\mathbf{X}\|^{2}}\right)\mathbf{X}$
@@
Assume that $\mathbb{P}_{\mathbf{X}\times Y}$ is defined on a countable domain $\mathcal{X}\times\mathcal{Y}$ with $\mathbf{0}_{n\times d}\notin\mathcal{X}$. Assume that parameter matrices $\{K_{i},Q_{i}\}_{i\geq 1}$ in the victim model $f$ have uniform bounded norms, i.e., $\|K_{i}\|\leq D$ and $\|Q_{i}\|\leq D$ for some $D>0$. There exists an $\alpha^{*}\in(0,1)$ depending on DD such that the following two statements are true.
(1) If $\alpha<\alpha^{*}$ and $\{K_{i},Q_{i}\}_{i\geq 1}$ are parameter matrices of the victim model, with $\hat{K}_{\alpha L}$ and $\hat{Q}_{\alpha L}$ as distilled parameters drawn from a continuous distribution on $\mathbb{R}^{n\times d}$, the matrix $\hat{f}_{\infty}(\mathbf{X})$ almost surely has rank one for all inputs $\mathbf{X}$.
(2) If $\alpha>\alpha^{*}$, there exists a victim model with parameter sequence $\{K_{i},Q_{i}\}_{i\geq 1}$ such that for any distilled parameters $\hat{K}_{\alpha L}$ and $\hat{Q}_{\alpha L}$, the matrix $\hat{f}_{\infty}(\mathbf{X})$ has rank greater than one for some $\mathbf{X}$.
Remark 1: The proof is provided in Appendix A. This theorem demonstrates that if the distilled parameters of the bottom layers (i.e., $\alpha<\alpha^{*}$) are obtained through a randomized algorithm, such as stochastic gradient descent, with a continuous distribution supported on $\mathbb{R}^{n\times d}$, the distillation will certainly fail, as the feature matrix degenerate. In contrast, keeping the later layers secured (i.e., $\alpha>\alpha^{*}$) does not maintain this property, indicating that it is more effective to secure the bottom layers before the transition layer, rather than the later ones. Further remarks are in Appendix A.5.
@@ 3.2 SOLID: Semi-Open Local Infrastructure Deployment Framework
Theorem 1 shows that securing bottom layers improves security. Inspired by this insight, we propose a method to approximately find the smallest bottom layer index set $I$ that satisfies $R(I)\leq(1+\varepsilon)R([L])$ for any small $\varepsilon>0$. To achieve this, a straightforward implementation is to begin with A simple approach is to start with $I_{l}=\{1,\ldots,l\}$ for each $l$ beginning from 1, then evaluate the distillation ratio $R(I_{l})$ after the attack, and identify the smallest $l$ that meets the inequality. This extensive fine-tuning process is time-consuming, prompting the critical question: Can we create a fine-tuning-free metric that predicts LLM performance under model distillation attacks? Hence, our goal is to establish a metric directly correlated with the distillation ratio.
@@
Previous research (Choi et al., 2024; Bailly et al., 2022) indicates that the difference $\|\bm{\theta}_{\text{dist}}(I,\mathcal{D})-\bm{\theta}_{0}(I)\|_{2}$ is minimal in large networks compared to the dataset size $|\mathcal{D}|$. In models such as single-layer ReLU networks (Anthony et al., 1999; Zou et al., 2020), this difference scales as $\mathcal{O}\left(\frac{|\mathcal{D}|}{\sqrt{N}}\right)$ (Jacot et al., 2018; Wei et al., 2019), where $N$, the number of model parameters, far exceeds the dataset size in large language models (LLMs) (Dubey et al., 2024; Liu et al., 2024). The first term, independent of fine-tuning, dominates and effectively predicts the distillation ratio. We refer to this term as the Distillation Difficulty (DD(I)(I)), defined as
@@
[…]